# Carroll black holes in (A)dS and their higher-derivative modifications

### v2

@@ III.2 Thermodynamics
If $10<9\Lambda M^{2}<$ 1, there are two horizons in the Lorentzian case which translates into two extremal surfaces in the Carrollian case, namely, a black hole extremal surface with entropy $\mathscr{S}_{H}$ equal to (58), and a cosmological extremal surface with entropy $\mathscr{S}_{C}=\tfrac{A_{c}k}{4cG_{M}\hbar}$, where $A_{c}$ is the area of the cosmological extremal surface. The total entropy is
@@
If $10<9\Lambda M^{2}<$ 1, i.e, we have a cosmological extremal surface, its surface gravity and the temperature would be similar to the black hole’s formula with $r_{H}$ is replaced by the radius of the cosmological extremal surface $r_{c}$, however, we have an effective temperature of [68, 69, 70]
@@
which also agrees with the result from Smarr-Gibbs-Duhem formula in [64]. If $10<9\Lambda M^{2}<$ 1, the cosmological extremal surface will have energy given by a formula similar to $A_{H}$ replaced by $A_{c}$, and the total energy would be the sum of the two. Notice that the energy in this case also goes to zero faster than in the case of Carroll Schwarzschild and Carroll Schwarzschild-Bach black holes.
@@ IV.1 Geodesics
×$\times[1-\tfrac{\Lambda}{u_{0}^{2}}+\delta(3-\tfrac{\Lambda}{u_{0}^{2}}-\tfrac{3}{8u_{0}^{2}k})+2\delta^{2}]\bigg{(}b+\frac{1}{u}\bigg{)}\sqrt{1-\frac{1}{bu}}F\bigg{(}\sin^{-1}\bigg{(}\frac{\sqrt{\frac{b+\frac{1}{u}}{b}}}{\sqrt{2}}\bigg{)}|\frac{2bu_{0}}{bu+1}\bigg{)}\bigg{)}\bigg{)}.$
Similar to the results in Sec. II.1, the deflection angle diverges at $u=1/b$, so we introduce the regularization parameter $a$ and evaluate it at $u=1/b-a$. The difference in this case is that $f_{1}$ and $f_{2}$ can be infinite when taking $\Lambda\to\infty$. For any finite value for $\Lambda$, the deflection angle diverges as in Sec. II.1, and we come to the same conclusion that adding higher-derivative terms implies that particles close enough to the extremal surface will wind an infinite number of times. In the limit $\Lambda\to\infty$ such that it grows faster than $a^{1/3}$, then $\phi$ goes to zero, however, such spacetimes cannot accommodate black holes [72], thus we neglect this possibility.
@@ IV.2 Thermodynamics
where $\mathscr{S}^{(1)}$ is the leading order, $\mathscr{S}^{(2)}$ is the next-to-leading order and $\mathscr{S}^{(3)}$ is the next-to-next-to-leading order with all higher orders are zero. In the case when $10<9\Lambda M^{2}<$ 1, we have two extremal surfaces as discussed before and the total entropy is the sum of the two entropies. In the case of (4,2), the entropy is
@@
If $10<9\Lambda M^{2}<$ 1, then we have a cosmological extremal surface and the energy for it is given by similar formulae, but replacing $r_{H}$ and $A_{H}$ with $r_{c}$ and $A_{c}$ respectively. The total energy is the sum of both extremal surfaces’ energies. The energy for the (4,2) case is given by
@@ V Conclusions and outlook
In addition to geodesic particle dynamics, we analyzed the thermodynamics of such black holes. The most notable aspect is the divergence of entropy when the temperature goes to zero, which is also present in Carroll Schwarzschild’s case. This means that there are an infinite number of microstates to these black holes, each of which has an infinitesimal probability. This agrees with the fact that the specific heat is calculated to be divergent, i.e. since the black holes have an infinite number of microstates, they can accommodate infinite thermal energy in a finite volume. Similar to the Lorentzian case, this leads to a classification of black holes according to whether the specific heat is positive or negative; this classification is harder in Schwarzschild-Bach-(A)dS due to the presence of an additional parameter (the Bach parameter). However, a notable simplification from the Lorentzian case is that Carroll black holes can be considered as incompressible thermodynamical systems. This arises from the facts that particles can not enter the extremal surface, and Carroll black holes do not radiate. Thus, the specific heat at constant pressure and at constant volume are the same, and the classification is simplified from the Lorentzian case, where we have to take all variables into consideration.

### v3

@@ III.2 Thermodynamics
If $0<9\Lambda M^{2}<1$, there are two horizons in the Lorentzian case which translates into two extremal surfaces in the Carrollian case, namely, a black hole extremal surface with entropy $\mathscr{S}_{H}$ equal to (59), and a cosmological extremal surface with entropy $\mathscr{S}_{C}=\tfrac{A_{c}k}{4cG_{M}\hbar}$, where $A_{c}$ is the area of the cosmological extremal surface. The total entropy is
@@
If $0<9\Lambda M^{2}<1$, i.e, we have a cosmological extremal surface, its surface gravity and the temperature would be similar to the black hole’s formula with $r_{H}$ is replaced by the radius of the cosmological extremal surface $r_{c}$, however, we have an effective temperature of [76, 77, 78]
@@
which also agrees with the result from Smarr-Gibbs-Duhem formula in [72]. If $0<9\Lambda M^{2}<1$, the cosmological extremal surface will have energy given by a formula similar to $A_{H}$ replaced by $A_{c}$, and the total energy would be the sum of the two. Notice that the energy in this case also goes to zero faster than in the case of Carroll Schwarzschild and Carroll Schwarzschild-Bach black holes.
@@ IV.1 Geodesics
$\times[1-\tfrac{\Lambda}{u_{0}^{2}}+\delta(3-\tfrac{\Lambda}{u_{0}^{2}}-\tfrac{3}{8u_{0}^{2}k})+2\delta^{2}]\bigg(b+\frac{1}{u}\bigg)\sqrt{1-\frac{1}{bu}}F\bigg(\sin^{-1}\bigg(\frac{\sqrt{\frac{b+\frac{1}{u}}{b}}}{\sqrt{2}}\bigg)|\frac{2bu_{0}}{bu+1}\bigg)\bigg)\bigg).$
Similar to the results in Sec. II.1, the deflection angle diverges at $u=1/b$, so we introduce the regularization parameter $a$ and evaluate it at u=1/bau=1/b-a. The difference in this case is that $f_{1}$ and $f_{2}$ can be infinite when taking $\Lambda\to\infty$. For any finite value for $\Lambda$, the deflection angle diverges as in Sec. II.1, and we come to the same conclusion that adding higher-derivative terms implies that particles close enough to the extremal surface will wind an infinite number of times. In the limit $\Lambda\to\infty$ such that it grows faster than $a^{1/3}$, then $\phi$ goes to zero, however, such spacetimes cannot accommodate black holes [80], thus we neglect this possibility.
@@ IV.2 Thermodynamics
where $\mathscr{S}^{(1)}$ is the leading order, $\mathscr{S}^{(2)}$ is the next-to-leading order and $\mathscr{S}^{(3)}$ is the next-to-next-to-leading order with all higher orders are zero. In the case when $0<9\Lambda M^{2}<1$, we have two extremal surfaces as discussed before and the total entropy is the sum of the two entropies. In the case of (4,2), the entropy is
@@
If $0<9\Lambda M^{2}<1$, then we have a cosmological extremal surface and the energy for it is given by similar formulae, but replacing $r_{H}$ and $A_{H}$ with $r_{c}$ and $A_{c}$ respectively. The total energy is the sum of both extremal surfaces’ energies. The energy for the (4,2) case is given by
@@ V Conclusions and outlook
In addition to geodesic particle dynamics, we analyzed the thermodynamics of Carroll black holes. The most notable aspect is the divergence of entropy when the temperature goes to zero, which is also present in Carroll Schwarzschild’s case. This means that there are an infinite number of microstates to these black holes, each of which has an infinitesimal probability. This agrees with the fact that the specific heat is calculated to be divergent, i.e. since the black holes have an infinite number of microstates, they can accommodate infinite thermal energy in a finite volume. Similar to the Lorentzian case, this leads to a classification of black holes according to whether the specific heat is positive or negative; this classification is harder in Schwarzschild-Bach-(A)dS due to the presence of an additional parameter (the Bach parameter). However, a notable simplification from the Lorentzian case is that Carroll black holes can be considered as incompressible thermodynamical systems. This arises from the facts that particles can not enter the extremal surface, and Carroll black holes do not radiate. Thus, the specific heat at constant pressure and at constant volume are the same, and the classification is simplified from the Lorentzian case, where we have to take all variables into consideration.